OpenStudy (anonymous):

Simplify the expression.-5+i/2i ,please. :)

OpenStudy (anonymous):

also explain

OpenStudy (amoodarya):

\[\frac{-5+i}{2i}=\\\frac{-5+i}{2i}*\frac{i}{i}=\\\frac{-5i+i^2}{2i^2}=\\\frac{-5i+(-1)}{2(-1)}=\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks